Draw the following network and calculate all dates (all activities are contiguous):

Activity

Duration

IPA

Type Rel

Lag

ES

EF

LS

LF

TF

FF

A

4

-

B

3

-

C

8

A

D

7

A,B

E

6

D

SS

3

F

3

D

FF

G

5

F

SS

2

2

F

FF

please do not copy from others, they are wrong
Answer #1

Answer:

A ! F D (9) 3 19 SS: 3

TF=LF-EF

=LS-ES

FF=MIN(ES of sucesors)-EFof activity

Activity ES EF LS    LF    TF    FF   
A 0 4 0 4 0 0
B 0 3 1 4 1 1
C 4 12 11 19 7 7
D 4 11 4 11 0 0
E 11 17 13 19 2 0
F 11 14 11 14 0 0
G 14 19 14 19 0 0